You want to buy a new sports car from Muscle Motors for $58,000. The contract is in the form of an annuity due for 36 months at an APR of 8.25 percent. What will your monthly payment be?
Answer #1

The present value of an annuity due = Monthly payment + Monthly Payemnt * [ {1 - 1/(1 +rate) ^ (period -1 ) } / rate]

Present value = $58000

Monthly payment = P

rate = APR/12 = 8.25%/12 =

period = 36

58000 = P + P * (   {1 - 1/(1+ .6875%) ^35)} / .6875%) =

58000 = P + P * 31.01325

P = 58000/32.01325 = $1811.75